TCCON > Sites > Reunion Island > Instrument History

Instrument History

Table of contents
  1. 1.  2017
  2. 2.  2016
  3. 3. 2015
  4. 4. 2014
  5. 5. 2013
  6. 6. 2012
  7. 7. 2011
  8. 8. 2010


  • September 25, 2017
    The instrument is working on the automatic mode and we start with out automatic solar measurement again.
  • September 22, 2017
    New computer is in place at St. Denis. All software are in place for the operation. Christian tried to do some cell measurements manually.
  • September 12 - 21, 2017
    Memory problem with the computer
  • September 4, 2017
    Change of the acquisition card
  • August 27, 2017
    No more connection with the instrument, instrument down
  • August 25, 2017
    Last date of measurements with the timing error problem. This problem started on July 4. The computer was not so stable during this time. One of the hard disk became unavailable for some time. The hard disk came back on its own but some data was lost. Unfortunately the old log file from the NetTime programme was also lost and the new file is written starting from a new date after August 25. This makes the analysis of the timing error very difficult.
  • August 2, 2017
    Restart of measurements
  • July 14 - 19, 2017
    Instrument down due to electrical maintenance in the building. 
  • July 4, 2017
    Start of the timing error problem for the measured atmospheric data.
  • June 2 - 15, 2017
    Problem with the Suntracker being stuck on the switch again and again several times during this period. Jean-Marc helped to unblock the suntracker but it got stuck again after some time. The problem seems to be similar to the one we had in May. Finally the problem with the suntracker is solved on 15 May and we performed some solar measurements. During the 2 June and 15 June period we continued to perform some lamp measurements.
  • May 10-12, 2017
    Problem with the suntracker being stuck. No measurement can be performed. It was repaired on 12 May. We start performing solar measurements again on 13 May.
  • April 18, 2017
    Jean-Marc communicated that the power is back at the station. We start our TCCON measurements again this afternoon.
  • April 9 - 17, 2017
    The system is down due to power cut at the station. The fault in the line is not detected.
  • March 15, 2017
    The power at the station is back. We start the measurements starting tomorrow.
  • March 10, 2017
    The system is down due to power cut at the station.
  • March 8, 2017
    The connection between the controller and the motors has been re-established. Our instrument is operational.
  • February 14, 2017
    The system is not operational due to XPS problem.
  • February 10, 2017
    We are re-launching the measures at 4 pm
  • February 7, 2017
    Instrument shut down due to a cyclone event occuring at Ile de La Réunion.
  • January 23, 2017
    Power is back at the station. Jean-Marc restarted the FTIR again.
  • January 22, 2017
    OCO-2 targetting for our site. Later in the evening power down due to maintenance works on January 23, 2017.
  • January 15, 2017
    OCO-2 targeting for our site.


  • December 21 - 22, 2016
    Electrical maintenance at the site. After the maintenance the UPS system did not start. So the FTIR is connected to the normal power line.
  • December 16, 2016
    OCO-2 targeting for our site.
  • November 29, 2016
    In the morning we went to the roof to un-mount the solar tracker (Jean-Marc and I). To our surprise we found that the solar tracker was in the correct home position. We decided to launch the auto-start program in BARCOS. Everything works fine. The solar tracker was at the right position and we started to perform TCCON measurements again. We will wait to see if this problem occurs again in the future. It may be related to the humidity. And when it rains on the day before the zero position sensor of the Newport doesn't work properly. We need to investigate this further.
  • November 28, 2016
    Christian contacted Newport to find a solution to the solar tracker error. Newport could not find the zero position of the solar tracker and thinks that the zero switch of the motor is dead. They need to have the electronic board of the azimuth motor to check on the problem. So the solution is to un-mount the solar tracker and send the electronic board to Newport for investigation. Christian closed all programs except the metro to get the meteorological parameters.
  • November 25, 2016
    Solar tracker error -5. The initialization of the azimuth motor of the solar tracker does not work anymore.
  • November 24 - December 4, 2016
    Site visit by Mahesh. TCCON HR and LR measurements performed.
  • September 9, 2016
    OCO-2 targeting for our site.
  • September 8, 2016
    UPS is reconnected and the system is back to operation with the UPS.
  • September 3-5, 2016
    Suntracker problem, no solar measurements performed. Only cell measurements performed on Sep 3.
  • August 26, 2016
    OCO-2 targeting for our site.
  • August 12, 2016
    UPS is disconnected and the instrument is connected directly to the power. The instrument is restarted and everything works fine.
  • August 4, 2016
    Power cut at the site. UPS drained out completely. The UPS has a failure and cannot restart. We need to wait till a solution is found. The instrument is non-operational until then. It is currently holiday time at the University.
  • July 12, 2016
    The instrument is operational again after 50 days. The construction works in not completely finished on the roof. They still need to do the side railing to secure the boundaries. But this work will not disturb our measurements, so Jean-Marc installed the suntracker and the meteo-station. I changed the aperture settings back to 0.5 mm (standard value). Nicholas did the evacuation of the Bruker for about 3 hours before starting the spectrometer again. This is needed for the start to avoid pressure fluctuation in the beginning. Everything seems to work fine.
  • May 23, 2016
    The instrument is shut down on the evening of May 22. Starting from today onwards there will be maintenance works undergoing on the terrace. This is necessary in order to prevent water infiltration. The estimated duration for the maintainace work is about 6 weeks. As a result we cannot perform any measurements during this time. Jean-Marc (thanks!) has covered the sun tracker to protect it from any damage.
  • May 11, 2016
    OCO-2 targeting for our site.
  • April 30, 2016
    Alignment of 4 diode of the solar stacker system by Christian.
  • April 24, 2016
    Change of Elevation mirror by Christian.
  • April 20, 2016
    OCO-2 targeting for our site.
  • April 15, 2016
    The suntracker is working again. On April 13 and April 14 the suntracker worked for short time and we collected 9 spectra on each day. Rest of the time it was showing error code: -26.
  • April 11, 2016
    The suntracker is down with error code: -26. Jean-Marc is out-of-office until 25/04/2016, so no local support available until he is back.
  • March 11, 2016
    The suntracker motor voltage was increased again by Jean-Marc to make it work.
  • March 8-9, 2016
    We are having problems with the sun tracker again. The camtracker software showed error code: -26. Jean-Marc did the re-adjustment of the potentiometer of the elevation motor. This solved the problem again and the suntracker works now.
  • February 29, 2016
    The Vaisala PTB210 Class A digital Barometer is not operational since 2:40:39. We do also measure pressure with Vaisala Weather Transmitter WXT510. New Vaisala PTB210 will be ordered and replaced soon.
  • February 24, 2016
    The suntracker works again. We are performing TCCON measurements again. We colleced 100 TCCON spectra today.
  • February 8-23, 2016
    The suntracker had a problem. The camtracker software showed error code: -26. In between the hydraulic pump broke for opening the hatch broke down. Jean-Marc replaced it with a spare pump. Christian made a tele-meeting with Newport and Jean-Marc - the problem was located. The potentiometer of the elevation motor was not correctly set and had to be re-adjusted. Jean-Marc did the adjustment and now it works :) . Starting from tomorrow we can collect data again.
  • February 7, 2016
    The suntracker mirrors are bad again due to the sea-salt conditions at the measurment site giving a lor SNR for the measured data. Christian increaded the aperture to get more signal and still be in the linear zone for the InGaAs detector. We will see the improvement of the SNR in the data starting from tomorrow.
  • January 15, 2016
    OCO-2 targeting.


  • November 26, 2015
    OCO-2 targeting.
  • November 12, 2015
    OCO-2 targeting.
  • November 01, 2015
    OCO-2 targeting.
  • October 11, 2015
    OCO-2 targeting.
  • September 21 - 25, 2015
    Instrument down for electrical maintenance works in the lab.
  • September 02, 2015
    OCO-2 targeting.
  • August 01, 2015
    OCO-2 targeting.
  • June 25, 2015 Azimuth and elevation mirrors of the sun tracker exchanged with new mirrors. Significant improvement in the SNR.
  • May 20, 2015
    OCO-2 targeting.
  • May 18, 2015
    OCO-2 targeting.
  • May 9, 2015
    OCO-2 targeting.
  • Filip Desmet left the group at the end of April. I, Mahesh Kumar Sha, am the new member in the group taking over the work from Filip.
  • April 2, 2015
    OCO-2 targeting.
  • March 24, 2015
    OCO-2 targeting.
  • March 1, 2015
    OCO-2 targeting.
  • February 18, 2015
    OCO-2 targeting: An almost completely sunny day.
  • January 25, 2015
    Currently on-site in St. Denis, during the previous week, I replaced the NI 9205 module so we have our meterology data back.  Also optimized the optical alignment of the interferometer, and replaced the broken halogen lamp.  Our tracker mirrors have been attacked by sea-salt, we were able to clean off the salt-deposits, but the protective layer on top of the gold surface has decayed significantly.  This is not really a problem for the longer wavelengths in the IR, but it causes some problems for our tracker camera which operates around 900 nm, I had to increase the aperture on the camera-lens.
    The sun finally appeared today, and I was able to verify that the tracking system is working correctly, so I restarted the TCCON measurements.
  • January 5, 2015
    The National Instruments 9205 module in the DAQ is not recognised by the computer, this means that either the module itself is broken, or there is a bad contact between the module and the NI DAQ.  Jean-Marc reinserted the module in the DAQ, and tried a different slot, but the module still wasn't recognised, meaning that the module itself is broken.
    We have a site-visit planned during the last two weeks of January, we will install a replacement module at that time.


  • December 26, 2014
    Our system can't access the "9205" analog input module of the "NI-DAQ" which provides the analog inputs for, amongst other things, the meteorological data.  Probably one of the power supplies (5V or 24V) is broken or a fuse has blown.
  • December 5, 2014
    One of our UPSs malfunctioned and caused a power-cut today.  We have removed this UPS, all of our critical equipment is now on the single remaining UPS.
  • November 13, 2014
    We had a power-cut in the afternoon, the system restarted automatically once the power got back.
  • October 22, 2014
    OCO-2 targeting: it was a completely sunny day, with continuous TCCON measurements.
  • October 20, 2014
    Camera tracking works fine, however, the intensity of the sun disk on the aperture wheel is less than what we had before.  Even with the camera gain and exposure time set to their maximum values, the intensity isn't what it used to be.
    We will ask Jean-Marc to check if the mirrors are clean, and perhaps open the aperture of the camera lens a little bit more.
  • October 17, 2014
    We've decided to wait a little bit longer before installing the O-ring for the elevation mirror again, to evaluate if the tracker finally keeps working as it should.
    It also appears that some settings of the tracking-camera have been changed.  Unfortunately we haven't had the opportunity to optimise them again due to cloudy, rainy weather over the last few days.
    Christian has already reduced the intensity value for the edge detection of the solar disk.
  • October 15, 2014
    We're operational again.
  • October 14, 2014
    The changes that we made to the tracker system appear to be working.
    On Thursday, Jean-Marc will put the O-ring back, but will add some silicone grease to deal with the friction problem.
  • October 13, 2014
    Last Wednesday, Jean-Marc removed an O-ring from between the solar tracker mirror and the elevation motor.   It appears that the rubber ring was causing too much friction which resulted in the motor getting stuck at certain points (initially only on the way down, but more recently also before noon).
    The O-ring however, also held the mirror in its place after we cut power to the elevation rotation stage.  Now that we have removed this ring, cutting power to the motor, causes the mirror to "drop", facing the Zenith.  The extra influence of gravity (added acceleration) also resulted in the rotation stage not being able to find its home position.  We adapted the tracker software, reducing the speed of the homing sequence from 360°/s to 2°/s and now it is working fine.
    We hope to restart our measurements tomorrow morning.
  • September 19, 2014
    Our strategy to stop the measurements before noon, and bring the elevation mirror back down at high speed did not succeed.  To do this, we have to hand over control of the tracker from our camera tracking program to the commercial control software, which requires a reinitialization of the motors and at this point we get again an error -26.
    I have added a function to our tracking software to bring the elevation mirror home at high speed, so that we don't have to make the transition to the Newport web-interface anymore (thus avoiding powering down the rotation stage).  Later today Jean-Marc will unblock the motor again, so that we can test this new feature.
  • September 15, 2014
    Jean-Marc has physically unblocked the elevation mirror again, it should be ready to go for tomorrow.
  • September 12, 2014
    We will stop the measurements during the weekend, next week we will change our scripts to end the measurements just before noon.  When we return the elevation mirror to its zero position at high speed it doesn't get stuck, it's only during actual tracking, when the elevation motor moves slowly, that we have this problem.
  • September 11, 2014
    OCO-2 targeting: our solar tracker got stuck again (error -26) just after local noon, we were able to measure throughout the morning.
  • September 10, 2014
    The tracker got stuck in the afternoon and threw an error -26.  We needed an on-site intervention of Jean-Marc to get the tracker moving again.
    We have been having problems in the afternoon for some time now, but until now we have always been able to get the tracker moving again using just the software interface.
  • September 8, 2014
    Jean-Marc has changed the oil in the vacuum pump today.
  • July 8, 2014
    Jean-Marc opened the instrument today after the measurements to remove the InSb detector (dewar needs to be evacuated).  It seems that we also have a problem with the NIR tungsten-lamp.  Until we find a replacement light bulb we will use the globar for our linefit measurements.
    The elevation rotation stage is also causing problems again, it runs fine in the morning, but in the afternoon, when the mirror moves down again, we get an error each time after just a few minutes of operation.
  • July 1, 2014
    The engineer from Micro-Controle did some additional tests of the system this morning and has adjusted some of the parameters in the motion controller.
    We have restarted the measurements.
  • June 30, 2014
    Christian and an engineer from Newport (Micro-Controle) did some remote tests of the elevation mirror rotation stage of the tracker.  At the end of the day Jean-Marc adjusted the position of the rotation stage so that it is now correctly centred.  Afterwards we were able to move the mirror up and down without getting any error-messages.
  • June 12, 2014
    We have done some extra tests with the elevation rotation stage of our tracker.  It turns out that when we had the same problem with our tracker on Maïdo, it was because of a wire that had come loose from one of the connectors on the motor.
    With the RGV100BL at St. Denis there is no problem with the soldering on the connectors of the rotation stage, and when we connect the rotation stage directly to the XPS motion controller it works fine.  This means that there must be a problem with the cable between the XPS and the roof, or with one of the cables or connectors of the solar tracker.
    We ran some additional tests later during the day, and the rotation stage now works again with the actual tracker cables, so there might just have been a bad contact with one of the connectors.  However now we have a mechanical problem with the rotation stage (error -22 after just a minute), possibly it hasn't been properly centred when remounting it.
  • May 23, 2014
    The elevation rotation stage of the tracker has stopped working, and refuses to reinitialize, it keeps throwing an error -26.  The error status in the XPS says "Thermistor fault" and "Hard interpolator encoder error".  A check with the calibration programme for the encoder (to the extent that this was possible) shows an ellipse rather than a cricle.
    The most likely cause of this problem is the encoder board in the rotation stage.  We've had this problem already with our other tracker on the Maïdo observatory.
    Fixing it means sending the elevation rotation stage back to Newport in France.  It looks like we will be nonoperational for a few weeks.
  • May 4, 2014
    An automatic update had caused a shutdown of the computer.  Had to restart it using the APC.  Everything is ready to go again for tomorrow.
  • April 29, 2014
    Jean-Marc removed the protection from around the tracker and reconnected all the cables.  The system is ready to start measurements again tomorrow.
  • April 28, 2014
    The Roofing works still haven't finished.  There are apparently some problems with the company that is supposed to carry out these works.
    Since the work on the roof won't continue any time soon, we will restart our FTIR measurements tomorrow.
  • March 17, 2014
    The roofing works still haven't finished, so the site remains inoperational.
  • March 3, 2014
    Today we stopped the measurements due to roofing works on the building in which the FTIR spectrometer is housed.  Jean-Marc has disconnected all the cables of the meteo station and the tracker, and he has covered the tracker to prevent any damage to our equipment.
    The roofing works on the section of the roof where our equipment is installed will take about 1 week.
  • February 14, 2014
    Last night there was a power cut.  The measurements will restart tomorrow, since today is quite cloudy.
  • February 5 - 6, 2014
    There was a read/write conflict with the hatch-status file.  The error must have occured some time yesterday but we only noticed it today as the error dialog box was well hidden behind all the other windows.
  • January 13, 2014
    Jean-Marc has reinstalled the meteo mast at St. Denis, and we have restarted the measurements.
    The cyclone took out some power lines to the Maïdo observatory, now 10 days later power still hasn't been restored to the mountain site.
  • January 4, 2014
    None of the instruments in St. Denis was damaged by the cyclone.  We will restart the measurements on January 13.
    We have no news yet of our instruments in the Maïdo mountain observatory, the Meteo France station on the mountain measured wind gusts up to 169 km/h.
  • January 3, 2014
    Yesterday in the late afternoon cyclone Bejisa passed some 40 km to the West of Reunion.  Winds of up to 150 km/h were measured on the island.


  • December 31, 2013
    We stopped the measurements at St. Denis and Jean-Marc secured all the equipment for cyclone Bejisa.
  • December 30, 2013
    A tropical cyclone is forecast for January 2 - 3, 2014.  Today we stopped the measurements at the Maïdo mountain site and prepared the equipment for the cyclone.  Today or tomorrow, we will also stop the TCCON measurements at St. Denis.
  • December 11, 2013
    We are operational, however there is still a problem with the vacuum which deteriorates quickly after pumping.  So until we find the leak, only our early morning measurements will meet the TCCON criteria for the instrument pressure.
  • December 10, 2013
    Today there is a planned power cut for a large part of the University, including the FTIR lab, from 7:00 until 7:30 Reunion time.
    It also turned out that all the instruments in the lab in St. Denis: FTIR, Picarro and UV-vis spectrometer were connected to one and the same power outlet.  Jean-Marc connected all of the FTIR equipment to a different outlet with its own breaker.  We hope that this will put an end to the frequent unexpected power-cuts that we have had these last few months.
    Jean-Marc also looked for a leak in the vacuum of the Bruker 125 but couldn't find anything.  So we'll wait and see if the vacuum improves over the next few days after repeated pumping.
    There were no measurements today, but everything is set for tomorrow.
  • December 9, 2013
    Jean-Marc once again restored power to the instrument.  It is quite cloudy so we did not restart the measurements.  We powerd down all the equipment for the planned power cut tomorrow.
  • December 8, 2013
    Nicolas was unable to connect to the instrument today, it appears that we have another power cut.  No measurements today.
  • December 6, 2013
    Jean-Marc arrived back at the Island and reset the circuit breaker at 12:00 UT, so we finally have power again.  The system is ready to start measurements tomorrow.
    I corrected a bug in the software of our solar tracker, the theoretical positions of the tracker were not calculated correctly now that the Sun appears in the South at noon, and passes the blind spot of our tracker.
  • November 26, 2013
    We got confirmation that there has been a power-cut in the FTIR lab.  Our usual contact person is abroad at the moment, and the people at Reunion are awaiting his instructions on how to get all the instruments in the lab (it's not just our FTIR) going again so for now, the power is still off.
  • November 24 - 25, 2013
    We're unable to connect to the computer in St. Denis.  We also can't reach our cameras, so it must be a network issue.
  • November 22, 2013
    Since yesterday morning we haven't been able to connect to the computer in St. Denis.
    This afternoon we found out that all the equipment connected to the outlets of the APC, including the computer, had been switched off.  We're not sure how this happened, although the log shows that the APC was rebooted yesterday shortly after 13:00.
    Jean-Marc confirmed that there was a power-cut on November 21.  However this shouldn't change the settings of the APC.
  • November 20, 2013
    Jean-Marc removed the MCT detector (which is not being used, since we have a CaF2 beamsplitter for the TCCON measurements) from the St. Denis spectrometer to test the Maïdo instrument with this detector.  The MCT detector of the Bruker 125/HR on Maïdo is currently not working, and we want to check if it is a problem of the detector or the electronics.
    There will be no valid TCCON measurements at St. Denis today because we can't evacuate the Bruker without the MCT detector in place.
  • October 4, 2013 - Site visit
    We stopped the measurements this morning, we will not be operational for a few days while we upgrade some of the systems on our site.
    Today we took down the old solar tracker and installed a new tracker.  The new tracker is an updated version of the old one.  It features a lightweight plastic hatch and a high precision Newport rotation stage for the elevation motor.
  • October 3, 2013 - Site visit
    In the evening, after the last measurement, we replaced the PTB210 pressure sensor with a new one which was calibrated last week at the Royal Meteorological Institute.
  • September 16, 2013
    We had yet another power-cut during the morning, we lost about 2 hours of sunlight.
  • September 11, 2013
    The computer was off this morning, so there probably was another power cut last night.
  • September 9, 2013
    This morning there was a problem during the start of the measurements, a reset of the computer was needed to get everything going again.  This problem appears to be related to a communication problem with the XPS motion controller, it is most likely related to the inproper end of the measurement cycle yesterday.
  • September 8, 2013
    There was a problem with the automation of the measurements.  It is unclear what is the cause.
  • August 24, 2013
    Measurements stopped at 8:00 UT when disk space on the computer ran out.  Christian transferred data to the second disk to free up space.
  • August 14, 2013
    There was a brief power-cut this morning.
  • July 29 - August 1, 2013
    No measurements, we stopped the system for an announced power-cut.
  • July 16 - 17, 2013
    No measurements, we stopped the system for an announced power-cut.
  • July 11, 2013
    There was a power-cut this morning, Christian has restarted everything.
  • July 10, 2013
    There was a problem with the XPS motion controller this morning, Christian started the measurements around 6:30 UT.
  • July 5, 2013
    Christian started the measurements at 9:00 CET, after the announced power-cut earlier this morning.
  • July 4, 2013
    We got an email from Jean-Marc, saying that there will be a power-cut tomorrow morning between 7:00 and 7:30 Reunion time.
    Because of a planned inspection of the electrical systems at the University of Reunion, there will be power-cuts on July 16-17 (between 7:00 and 8:00 Reunion time), July 29 and July 31.  For these last two dates no time has been specified.
  • June 20, 2013
    An update about the oil level of the vacuum pump and the water level of the cooling unit from Jean-Marc:

    Maïdo :
    - niveau d'huile : plein, huile bien claire
    - niveau d'eau : plein

    StDenis :
    - niveau d'huile : 3/4, huile un peu opaque
    - niveau d'eau : 3/4

  • June 6, 2013
    There was another power-cut after 18:00 local time yesterday.  This morning Jean-Marc restored the power to the instruments.
  • June 5, 2013
    There have been two power cuts today, one during the night or early morning, a second one later on after 10:30 Reunion time, it's cloudy at the moment so we will skip the afternoon measurements.  No data today.
  • May 16, 2013
    There was a power-cut at St. Denis this morning, Christian restarted the system at 8:30 CET.
  • May 13, 2013
    Last week, we stopped the measurements on May 8 and 9, we also lost May 10 because the meteorological measurements didn't properly start.
    There currently are some problems with our liquid nitrogen supply at Reunion, this means that for now we can't run any NDACC measurements.
  • May 2, 2013
    Jean-Marc notified us that due to network maintenance at the University of Reunion, there will be no internet access on May 8 and 9.
    We will stop our measurements during these two days.
  • April 11, 2013
    There appears to have been a power-cut last night.  Nicolas restarted the automation and control software, there were some software problems with the controls for the evacuation of the Bruker and the hatch, but eventually he managed to get everything running.
  • April 2, 2013
    Jean-Marc has installed the replacement power supply, the optical reference switch for the elevation rotation stage is now working correctly.
    The measurements should restart tomorrow.
  • March 27, 2013
    We tracked the problem with the elevation motor's optical reference switch to a faulty power supply (5 V/24 V bivolt power supply).
    We have a spare power supply in Brussels which we will ship to St. Denis tomorrow.
  • March 25, 2013
    Together with Jean-Marc we tested the elevation motor of the solar tracker.  We were able to pinpoint the cause of the problems to the optical reference position switch of the elevation.
  • March 21, 2013
    There is a problem with the elevation motor of our sun tracker, it may just be stuck at its limit position, or it could be worse.  Jean-Marc will check later today or tomorrow.
  • March 20, 2013
    Last night a storm passed over Reunion Island, there was no damage at our site in St. Denis.
  • March 04, 2013
    The site is operational.
  • March 03, 2013
    I found the problem with the macros, OPUS 7 is more strict than OPUS 6.5 for the phase resolution of the spectra.  During an initial low-resolution test scan in each macro the phase resolution (PHR) is higher than the spectral resolution (RES) which causes OPUS 7 to throw an error.  I updated the macros to also send the phase resolution in the measurement commands, and I make sure that PHR = 2*RES.
  • February 18 - 19, 2013
    Short site visit by Christian and Gregor to install a new laser.   The acquisition software was upgraded from Bruker OPUS 6.5 to OPUS 7, and the instrument got a firmware update.  There is a problem with the macros for the automated measurements, so for the moment no measurements.
  • February 5, 2013
    The IT department at the University of Reunion has found the cause of our intermittent problems with the FTP transfers to Brussels.  The computer was configured with two different gateways (one to the internet, one to the local instrument LAN), so half of the time it tried to connect to the remote FTP server over the local LAN instead of over the internet.
  • January 9, 2013
    The voltages of the two laser detectors are now both 0V.  Meaning that the laser shutter won't even open anymore.
  • January 8, 2013
    The system hasn't been running since last Friday, Jan 4.  Whenever why try a measurement with a resolution higher than 2 cm-1 we get a laser error.
    We will install a new laser during a site visit planned for February 18-19.
  • January 2, 2013
    A cyclone is approaching the island, Jean-Marc has taken down the meteo mast and we have suspended the measurements.
    The cyclone should pass 100 - 200 km to the West of Reunion Island on the third of January.


  • December 20 - 31, 2012
    The laser errors are back again.  We enabled the "laser amplifier" option in the beamsplitter EEPROM (we are currently using the KBr BMS ID chip while we wait for a replacement CaF2 chip) but this doesn't seem to help.
    The current BMS ID chip is giving us problems with the name of the beamsplitter which is often returned as "KBrÿÿÿÿÿÿÿÿ", this causes problems with our OPUS macros.
    It also appears that OPUS is not saving all of the spectra that it records.
    The weather wasn't very good during this period so we didn't loose too much data.
  • December 12, 2012
    We have been having some problems with the calculation of the theoretical trajectory for our solar tracker since we lost the connection to the instrument last weekend.  Today we reverted to a previous version of the solar position algorithm (from Sept. 2011), which seems to be working fine.  The offsets also appear to have changed (not sure why) we'll see what happens during the next few days before updating them.
  • December 10, 2012
    We lost the internet connection to the instrument last Saturday evening (Dec. 8).  There is no problem with the network infrastructure at the University of La Réunion, so it may be our router which is causing these problems.
    We currently have limited on-site support since Jean-Marc is on vacation.
    Resolved: there was a problem with the network switch for the lab.
  • November 24, 2012
    We are operational again.
  • November 23, 2012
    The replacement beamsplitter ID card from Bruker still hasn't arrived at Réunion.  We have now replaced the defective BMS ID card of the CaF2 beamsplitter with the card from the KBr beamsplitter.  I have updated all the XPMs, and if all goes well, we will be operating again tomorrow.
  • November 21, 2012
    I installed a new version of the time sync service on the computer (NetTime version 3.14 - this new version keeps a log of the changes that it makes to the system time.
    In the meantime we are still waiting for the new beamsplitter ID card to arrive.
  • November 12, 2012
    Jean-Marc changed the address of the time server on the computer.  It has come to light that there is a problem with the time server at the University of Réunion that we have been using until now, the time returned by this server differes from UT by about two minutes (2 minutes late vs. UT).  Now we have put an NIST server as our preferred time server.
    At the moment we don't know yet for how long we have been using an incorrect time, we will try to find out as soon as possible.
  • November 8, 2012
    Christian contacted Gregor about our problem with the beamsplitter ID chip,

    Gregor writes:

    Dear Christian,
    The problem is well known. I guess the beamsplitter ID board is defective.
    The problem we have with these boards is that the chip with the three gold
    contacts is badly soldered onto the print. The problem is due to those
    solder points they are underneath the component which makes it rather
    difficult to solder them proper.
    Maybe somebody in an electronic workshop may help you. Or if you prefer we
    will send you a new print I24216 free of charge.

    You may check the beamsplitter ID chip :
    Access by : Browser-> Service-> Edit Hardware-> Edit Beamsplitter Eprom.

    Best regards,


    When we try to check the beamsplitter ID chip, we a "No beamsplitter EEPROM available" message.
    We will ask Bruker for a replacement print board.
  • November 7, 2012
    OPUS doesn't recongnise our beamsplitter anymore, instead of "CaF2" it returns some nonsensical string.  Usually we can solve this with by letting OPUS reconnect to the FTS, or by resetting the instrument, but now none of this is working.  Jean-Marc even opened the instrument and cleaned the chip on the  beamsplitter, but to no avail.
    At the moment we aren't running because all our macros expect a "CaF2" beamsplitter string to work.
  • September 25, 2012
    Today and last Sunday, the 23rd, our Newport XPS motion controller returned a following error, each time at about 12:15 UT.  Most likely something is causing a bit too much resistance for the elevation rotation stage (a loose cable or some mechanical wear).  A reinitialisation of the rotation stages gets the tracker going again.
  • September 18, 2012
    Since Friday the 14th of September we have been experiencing intermittent problems with our internet connection to our instrument.  We stopped operations during the weekend (16th and 17th) because of these problems.  Jean-Marc reported to us this morning that the problems with the connection to the University of La Réunion should now be fixed.
  • August 28, 2012
    This morning the tracker opened correctly, the laser hasn't caused any problems since we have restarted it last Friday.
    During the afternoon, Jean-Marc had a look at the tracker and found that one of the cables for the limit switches on the pneumatic cylinder was hanging in the path of the hatch, preventing it to open until the tracker had moved to a certain azimuth.
  • August 27, 2012
    The laser made it through the weekend without needing a restart.
    The hatch of our solar tracker is now also causing us problems.  It doesn't open completely until about local noon at La Réunion.  Tomorrow or Wednesday Jean-Marc will try to find the cause of this problem.  In the meantime we can get the hatch to open by moving the tracker to its mid-day position.
  • August 26, 2012
    I got an email from Debra with a temporary fix for our laser shutter problem.  However the cause of or problems is the fact that our laser is on its way out.  We're hoping that it will last until December when we have planned a site visit.
  • August 24, 2012
    It appears that we are having problems with our laser now.  The shutter underneath the beamsplitter compartment which lets the laser light into the interferometer opens and then closes again all the time.  Yesterday before we could run the test scans on our instrument, we had to turn the laser off and on again (at that time the error LED on the laser power supply was lit).
    We restarted the laser, and the IFS is now working again, and we are collecting data for TCCON!  This is however the second time in two days that we've had to restart the laser to get it working correctly.
  • August 23, 2012
    Jean-Marc has replaced the 4-way multiplexer card in the detector compartment of the IFS with a new card from Bruker.  Our initial tests show that with this new card, we can once again acquire good interferograms.
    If all of our tests go well today, then we should be operational again tomorrow.
  • July 19, 2012
    Today Jean-Marc tested our instrument by directly connecting the InGaAs detector to the IDET cable, thus bypassing the multiplexer card in the detector compartment of the Bruker 125.  In this configuration we have no artifacts in our interferograms.  So the problem that we are seeing must be related to either the muliplexer card itself or one of the flat cables connected to it.

    The image shows the InGaAs detector connected to the IDET cable in the HCl cell compartment.
  • July 18, 2012
    We have been in contact with Bruker about the problem with our FTS.

    Gregor writes:

    The problem we face here seems not just a simple laser disalignment. Please
    try wethere all available detectors show this type of damaged IFGs. When you
    run "Check Signal" does it also show a damaged interferogram?
    Is looks like the digitized data flow is currupted either a the detector
    unit or at the four way multiplexer inside the detector compartment.
    Also the IDET link cable to the SCT board or the SCT board it self may cause
    the problem.

    All our detectors do indeed display interferograms, with sudden jumps in the detector voltage.

    Tomorrow we will check if the four way multiplexer in the detector compartment is to blame.  To check this we will bypass the multiplexer by removing the InGaAs detector from its position, and connecting it directly to the cable (IDET) which is connected to the flange under the detector compartment.  If we can then collect a decent interferogram with the InGaAs detector in this way, we know that the multiplexer is to blame, otherwise it is something further down the line.  See test_july_2012.pdf for some pictures.

    No matter what the outcome of this test will be, we will not be operational again until late August, since the Université de La Réunion where our instrument is based will be closed from July 20 until August 12.
  • June 28, 2012
    Yesterday, Christian contacted Gregor about the problem with our Bruker 125/HR.
    Gregor suggested to change the gain of the laser amplifier.
    In the web interface of the Bruker 125: Service > Edit hardware configuration > Edit Beamsplitter EEPROM > check the box next to "Laser Amplifier On" > Click "Save".  (This amplifies the laser signal by a factor of 3)
    However, this didn't solve our problem.
    Today I took a closer look at some of our interferograms (img-ifgs-stdenis.pdf), there are regular sudden jumps in the detector voltage, so we may have a problem with our ADC or acquisition hardware.
  • June 22, 2012
    We haven't been able to measure a decent interferogram since June 18.  There appears to be some laser-related problem, however during our measurements OPUS doesn't give an error (as it normally does when the laser modulation becomes too small), yet the collected spectra are complete rubbish.  The strange thing is that when we send the mirror to the back of the instrument we almost instantly get an error that the laser modulation is too low.
  • June 18, 2012
    Last week (during the Wengen IRWG-TCCON conference), we started experiencing some problems with our high-resolution NDACC scans, most of the collected NDACC spectra are bad, the TCCON spectra however appear to be not affected by the problem.  This seems to indicate that there is a problem with our laser (or its power supply) or the alignment of the laser at high OPDs.
  • May 17, 2012
    We found the cause of our problems with the rain-detector.  Apparently, the internal configuration of the Blackbox RS-232 to ethernet converter, to which the rain detector is connected, got changed somehow.  We reconfigured it and now everything is working again.
  • May 14, 2012
    Jean-Marc has replaced the MAX233 RS-232 converter IC but the problem with our rain detector persists.  We will now check for problems further down the line.
  • May 07, 2012
    Jean-Marc has tested all the electronics of the rain detector but found nothing unusual.  So now we're thinking that there might be a problem with the microprocessor.   We've sent a replacement processor to La Réunion.
  • Apr 30, 2012
    We're having some trouble with our U-shaped rain detector (Lambrecht), probably a problem with our electronic system.  We have suspended operations since the signal from the rain detector is used to decide when to open and close the hatch of our tracker.
  • Feb 14, 2012
    The meteo mast is back up again after the cyclone.
  • Feb 12, 2012
    With cyclone Giovanna approaching, Jean-Marc took down the meteo mast and secured the tracker in the morning.  The cyclone passed some 200 km North of La Réunion.  But still it was sufficiently strong to throw around some empty containers at the airport, and wind speeds of 230 km/h were measured at the Maïdo.
  • Feb 03, 2012
    The system is up and running again.
  • Feb 02, 2012
    Using the interface connector built by Francis, our colleagues at La Réunion measured the voltages going to and from the azimuth rotation stage at Newport XPS controller and the rotation stage.  The voltage at the rotation stage was found to be too low (probably due to the quite long cable between the controller and the rotary stage).  We installed a small DC-DC convertor box supplied by Newport between the cable and the azimuth stage, this slightly increased the voltage of the home signal (by 0.2 V).  As a result the azimuth stage can now once again find its home position, although we clearly don't have a lot of room for error here, and some mechanical vibration or movement of the cables or the pins in the connectors could result in a drop of the voltage and cause another malfunction.
    During the last weeks, when the system was not operational, we also had a problem with the vacuum of the instrument.  The pressure would rise to more than 10 hPa in a day after being evacuated, but that problem has now also been solved.
  • Jan 19, 2012
    Possible causes of the problems with the azimuth rotation stage of the solar tracker could be: a bad contact in one of the connectors (home position signal not reaching the controller), or according to Newport, the fact that we use quite long cables between the rotation stage and the motion controller, they propose to add an extra card to our controller which can supply a higher voltage.
    At the moment we are working on a connector to allow us to measure the voltage at the rotation stage and the controller.
  • Jan 18, 2012
    The good news is that the bug in the software has been fixed, the bad news is that the azimuth rotation stage of the tracker failed to find its home position and cannot be used.
  • Jan 17, 2012
    A bug in our software which takes care of the daily scripts prevented the system from starting automatically this morning, we are working to get it fixed.
  • Jan 16, 2012
    Jean-Marc managed to get our tracker unstuck again using a number 3 Allen key.


  • Dec 23, 2011
    We have the same problem with our solar tracker as on Dec 7, 2011.  Our colleagues at La Réunion tried to get it unstuck from its end-of-travel position earlier today but failed.
    At the moment it appears that we won't be operational again until at least Jan 16, 2012.
  • Dec 16, 2011
    The system is operational again.
  • Dec 15, 2011
    The rotary stage of the solar tracker was manually unstuck earlier this week, and the cables were disconnected and reconnected again.  The tracker has been tested over the last few days and is working correctly again.
    Today Christian is checking the alignment of the Bruker 125/HR with linefit.
  • Dec 12, 2011
    The Bruker 120/M performed its last NDACC measurements in St. Denis today, and will return home to Brussels.  The Bruker 125/HR is now our only instrument at La Réunion, and when its solar tracker is working again, it will continue to collect data for TCCON and NDACC.  The repair of the solar tracker is planned for later this week.
  • Dec 11, 2011
    Christian arrived at La Réunion today, during the next week he will disassemble our Bruker 120/M NDACC observatory in the rooftop container, and fix the solar tracker of the Bruker 125/HR.
  • Dec 07, 2011
    Failure of the solar tracker.  The tracker moved past its azimuth end-of-travel position and is now stuck.  A manual on-site intervention is required to get it going again.
    The last time when we experienced the same problem it was due to some dirt on a limit-sensor in the azimuth rotation stage.
    Fortunately, a site visit is planned for next week.  During this visit, Christian will decommission the Bruker 120/M which has been operating in a container on the same site.
  • Sept 05 - Sept 15, 2011
    Installation of the homebuilt solar tracker and new meteorological station at St. Denis, and start of the measurements with the Bruker 125/HR.  The 125/HR takes the place of our 120/M machine which has been performing NDACC measurements on the island during various missions since 2002, and continuously since June 1, 2009.  The 120/M is now installed in a container on the roof of the building, and will continue to measure on the island until November 2011.
    When we returned to the site, about one year after the installation of the 125/HR, we found the vacuum in the machine still to be less than 30 hPa.


  • Sept 20 - Oct 1, 2010
    Installation of the Bruker 125/HR at the Université de La Réunion in St. Denis.  The instrument is not yet used, and NDACC measurements on the site continue to be performed by the 120/M machine.


Tag page

Files 2

FileSizeDateAttached by 
IFS problem june 2012 - Interferograms of HBr and HCl cells with lamp.
636.88 kB05:54, 28 Jun 2012desmetActions
Bypassing the multiplexer in the Bruker 125
1687.85 kB05:47, 18 Jul 2012desmetActions